5. Below is an income statement for Jean Simmons Company:Sales$300,000 Variable costs(150,000)Contribution margin$150,000 Fixed costs(100,000)Profit before taxes$ 50,000 1) What was the company's margin of safety in dollars?2) If the unit sales price for Jean Simmons’s sole product was $10, how many units would it have needed to sell to produce a profit of $40,000?
Answer #1

Answer :

1) Margin of safety = $1,00,000

3) units to be sold = 28,000 units

Explanation :

1) Margin of safety = sales - Break even sales

Break even sales = (fixed cost ÷ contribution) × sales

= (1,00,000 ÷ 1,50,000) × 3,00,000

= 2,00,000

Therefore ,

Margin of safety = 3,00,000 - 2,00,000 = $1,00,000

2) units required to be sold to achieve target profit :

= (Fixed cost+ target profit) ÷ unit contribution margin

= (1,00,000 + 40,000) ÷ (10 - 1,50,000/30,000 units)

= 1,40,000 ÷ 5

= 28,000 units